# Dynamic-SUPERB Phase-2: A Collaboratively Expanding Benchmark for Measuring the Capabilities of Spoken Language Models with 180 Tasks

### v1

@@ Abstract
Multimodal foundation models, such as Gemini and ChatGPT, have revolutionized human-machine interactions by seamlessly integrating various forms of data. Developing a universal spoken language model that comprehends a wide range of natural language instructions is critical for bridging communication gaps and facilitating more intuitive interactions. However, the absence of a comprehensive evaluation benchmark poses a significant challenge. We present Dynamic-SUPERB Phase-2, an open and evolving benchmark for the comprehensive evaluation of instruction-based universal speech models. Building upon the first generation, this second version incorporates 125 new tasks contributed collaboratively by the global research community, expanding the benchmark to a total of 180 tasks, making it the largest benchmark for speech and audio evaluation. While the first generation of Dynamic-SUPERB was limited to classification tasks, Dynamic-SUPERB Phase-2 broadens its evaluation capabilities by introducing a wide array of novel and diverse tasks, including regression and sequence generation, across speech, music, and environmental audio. Evaluation results indicate that none of the models performed well universally. SALMONN-13B excelled in English ASR, while WavLLM demonstrated high accuracy in emotion recognition, but current models still require further innovations to handle a broader range of tasks. We will soon open-source all task data and the evaluation pipeline.
@@ 1 Introduction
Despite significant research in universal speech models, evaluating them effectively and comprehensively remains a major challenge. In NLP, to evaluate text LLMs, some benchmarks consist of a larger number of tasks. CrossFit (Ye et al., 2021) includes 160 tasks, BIG-bench (Srivastava et al., 2022) comprises 204 tasks, and Natural-Instructions (Mishra et al., 2022; Wang et al., 2022) provides 1,616 tasks. For universal speech models, several benchmarks have been proposed but with a fixed and limited set of tasks (typically around a dozen) to evaluate specific capabilities of models (Yang et al., 2021; Dunbar et al., 2022; Maimon et al., 2024). This is insufficient for evaluating a universal model, as we aim to investigate whether such models can perform a broad range of tasks beyond the limited set of tasks included in these benchmarks. There is a strong demand for a benchmark that can evaluate these models across various aspects with a wide range of speech tasks, which led to the creation of Dynamic-SUPERB (Huang et al., 2024a).
@@
Dynamic-SUPERB is the first benchmark designed to evaluate universal instruction-based speech and audio models. As universal models advance, we dynamically expand the benchmark by adding new tasks to provide better guidance for researchers in developing their models. We have designed a well-structured pipeline that harnesses the collective efforts of the entire community to gather a diverse array of challenging, novel, and creative tasks. The first generation of Dynamic-SUPERB consists of 55 tasks, covering various aspects of speech (e.g., content, semantics, speakers, etc.), making it the speech benchmark with the most tasks. However, this is not sufficient to pave the way for developing universal speech models, especially given the complexity and richness of the information conveyed by spoken language.
@@
This paper presents the Dynamic-SUPERB Phase-2. We collected 91 new tasks with contributions from the global research community, allowing the benchmark to grow more than twice as large as its first iteration. They cover a broad range of types, and some of these tasks present novel challenges that have not been explored in any previous research. Besides, previous speech research has typically treated music and environmental audio as background noise to be ignored. However, these sounds contain rich information and share overlapping elements that complement each other. Thus, in Dynamic-SUPERB, we also introduced preliminary music and audio tasks from established music and audio benchmarks (Yuan et al., 2023; Turian et al., 2022). We also define the core tasks by reformulating the SUPERB (Yang et al., 2021) (speech), MARBLE (Yuan et al., 2023) (music), and HEAR (Turian et al., 2022) (audio) benchmarks to enable quick-round research experiments. One challenge of evaluating with such a large-scale benchmark is deriving concrete and useful insights from hundreds of evaluation results. We provide a taxonomy for every task in Dynamic-SUPERB, where tasks are clustered by the specific model capabilities they probe. Researchers can follow this taxonomy to develop or reinforce specific capabilities of the models they build. To our knowledge, Dynamic-SUPERB is the largest benchmark and the first to provide such a detailed task taxonomy in speech processing.
@@
We conducted a comprehensive evaluation of several models using Dynamic-SUPERB Phase-2. To evaluate the diverse output formats of these models, we propose an automated pipeline that uses LLMs to assess and process model outputs for general evaluation across various tasks. The evaluation results show that these models perform well only on a very limited range of tasks. For example, SALMONN-13B performed well on English ASR, while WavLLM achieved high accuracy in emotion recognition. However, they do not generalize well to a much broader range of tasks in speech recognition and paralinguistics. On the other hand, the results reveal that training on diverse data, even with significant differences in signal-level characteristics, can enhance performance across domains. We observed that spoken language models outperformed music language models in certain music tasks. This highlights the potential for developing unified models for speech, music, and general audio processing. Importantly, we have open-sourced all materials, including task data and the evaluation pipeline, to facilitate reproduction and further development by researchers. We also invite researchers to join us in continuing to enhance Dynamic-SUPERB, making it more diverse and comprehensive to advance the development of universal spoken language models.
@@ 2.1 Instruction-following universal speech models
LLMs have shown strong natural language processing abilities and are used in speech, audio, and music applications. Recent frameworks integrate a pre-trained speech encoder with an LLM through fine-tuning techniques such as LoRA (Hu et al., 2021). LTU-AS (Gong et al., 2023) combines Whisper (Radford et al., 2023), an ASR system, with LLaMA (Touvron et al., 2023a), and it is fine-tuned on an open-ended speech and audio question-answering dataset. SALMONN adopts a window-level Q-former (Li et al., 2023a) to generate soft prompts that fuse representations from the speech and audio encoders. Qwen-Audio (Chu et al., 2023) introduces task-specific tags into Qwen to encourage knowledge sharing and prevent interference across diverse tasks, and it supports multi-turn dialogues for both audio and text inputs. WavLLM (Hu et al., 2024) uses curriculum learning to prevent overfitting on specific tasks while maintaining the LLM’s original capabilities. All these models accept speech, audio, and text as input but output only text, focusing on understanding rather than generation in speech and audio. There are also several attempts to use prompts for speech, music, and audio generation (Guo et al., 2023; Agostinelli et al., 2023; Liu et al., 2023a). However, to our knowledge, there is no universal model that is capable of handling both generation and understanding tasks. Moreover, these models have not been comprehensively evaluated on a benchmark, which hinders fair comparisons between them.
@@ 2.2 Evaluation benchmarks
Several benchmarks have been developed to evaluate the performance of speech models. SUPERB (Yang et al., 2021) is the most widely used benchmark for assessing the performance of speech foundation models across various tasks that cover different aspects of speech. On the other hand, SLUE (Shon et al., 2022; 2023) focuses more specifically on spoken language understanding. Yet, they are primarily limited to English. LeBenchmark (Evain et al., 2021; Parcollet et al., 2023) evaluates speech foundation models specifically for French, while IndicSUPERB (Javed et al., 2023) is dedicated to languages spoken in India. ML-SUPERB (Shi et al., 2023; 2024) offers ASR and language identification tasks in 100+ languages. Aside from speech, MARBLE (Yuan et al., 2023) and HEAR (Turian et al., 2022) offer platforms for evaluating various music and a wide range of audio tasks. Using these benchmarks, researchers build a specialized model for each task. Thus, the number of tasks is limited due to the growing costs associated with adding more tasks. Conversely, universal models are expected to do various tasks without fine-tuning for each one, allowing users to engage in far more diverse applications and enabling benchmark developers to include more tasks for comprehensive evaluation. In NLP and CV, benchmarks have been developed to evaluate models across a much broader range of tasks (Bitton et al., 2023; Srivastava et al., 2022; Mishra et al., 2022). However, in speech, we lack a benchmark of comparable scale. Dynamic-SUPERB (Huang et al., 2024a) is the first benchmark for instruction-based universal speech models. While its first version includes far more tasks than all other speech benchmarks, they are all classification tasks. AIR-bench (Yang et al., 2024b), on the other hand, includes tasks from speech, music, and audio, and expands beyond classification, although the number of tasks available for evaluating universal models remains limited (19 tasks in foundation track). Thus, there remains a critical need for a comprehensive benchmark that evaluates universal models across a broader range of tasks to fully assess their capabilities. Consequently, we developed Dynamic-SUPERB Phase-2, substantially upgrading the first generation with a detailed taxonomy and establishing community contribution protocols to facilitate continual task integration.
@@ 3.2 Task formulation
As Figure 1c depicts, in Dynamic-SUPERB Phasse-2, each task is structured to include: (1) Text instruction: A natural language instruction that guides the model on the task to perform. Each task has multiple different instructions to evaluate the model’s ability to understand instructions. (2) Audio component: At least one audio element, which can be in the input, the output, or both. (3) (Optional) Text component: Text elements (other than instruction) that may serve as inputs or outputs. The number of audio elements in the inputs or outputs may vary depending on the task. For example, in speaker verification, two utterances are used to determine whether they were produced by the same speaker (green blocks in Figure 1c). Besides, text inputs are not always required; for instance, ASR does not involve any text inputs. We use text format for instructions instead of spoken format. Spoken instructions involve varying content, speaker characteristics, and prosody, making them more complex. Text instructions act as an intermediary, bridging text and spoken universal models. These designs ensure consistency and simplify the model’s understanding and processing of diverse tasks while maintaining the benchmark’s extendibility with minimal constraints.
@@ 3.3 Call for tasks
Dynamic-SUPERB fosters community collaboration, encouraging the addition of innovative tasks to remain relevant in this rapidly changing field. In Phase-2, we initiated a call for tasks in March 2024 to invite contributions from the research community. We established an organized and transparent submission process on our GitHub portal, where we organize members who serve as editors to guide contributors. Contributors propose tasks by providing relevant information on GitHub. Each proposal is assigned to an editor who checks for major issues and prevents duplicated efforts. Then, task proposers upload their data to our Huggingface space in specified formats, using only datasets with proper licenses. They complete submissions by opening a pull request on GitHub with the necessary files. Afterward, editors review submissions on a rolling basis, offering suggestions for refinement rather than immediate acceptance or rejection. After iterative improvements, accepted tasks are merged into the repository. Between March and July 2024, we received about 145 task proposals and accepted 91 tasks. Tasks still under review are not included here but will be featured in the next phase. For details, please refer to Appendix F.
@@ 3.4 Task taxonomy
Figures 2 illustrates the high-level task taxonomy in Dynamic-SUPERB Phase-2. Due to the space limitation, we only show some representative tasks. Each leaf node includes at least one task and may be further categorized into more fine-grained sub-domains, which are not shown here due to space constraints. For example, within ‘Speaker & Language/Speaker’, we have two categories: ‘Speaker Characteristics’ and ‘Speaker Identification’, each containing several tasks. Please refer to Appendix B for the complete list of tasks. We first categorize tasks into two primary fields: (I) speech and (II) music & audio, which are generally distinguished by the source of the sound. For instance, speech is produced by humans, music is created using instruments, and audio includes sounds generated by other creatures, materials, or natural phenomena. We then split each field into several domains based on the key attributes and challenges that the tasks within them present.
@@ 3.4.1 Speech
As Figure 2(a) shows, there are 8 domains within the speech field. Speech Recognition focuses on converting spoken language into text. This includes tackling various challenges like multilingual and code-switching ASR, as well as spontaneous ASR which is very different from audiobook-style ones. It also contains specialized tasks such as command recognition and keyword spotting. Speaker and Language addresses the analysis of speaker characteristics and languages, covering tasks such as speaker verification, diarization, and language identification. Spoken Language Understanding deals with understanding and analysis of the content and semantics of spoken language. It covers tasks like sentiment analysis and speech translation. Phonetics, Phonology, and Prosody focuses on the sound structure of speech, including phoneme recognition, pronunciation evaluation, and prosodic features like stress and accent classification. Paralinguistics explores non-verbal aspects of speech, such as emotion recognition and vocal event detection, which can capture nuances like screaming or coughing. Speech Enhancement aims to improve speech quality by detecting and mitigating noise, reverberation, and other degradations. Speech, Voice, and Hearing Disorders is dedicated to identifying and classifying disorders such as stuttering and so on. Safety and Security focuses on detecting synthetic or manipulated speech, addressing tasks like spoof detection and recognizing deepfake voices.
@@ 3.4.3 Core tasks
While our task taxonomy provides a comprehensive, hierarchical, and systematic approach for benchmark users to easily get started with Dynamic-SUPERB Phase-2, evaluating all tasks with limited resources remains a significant challenge. To address this, alongside the task taxonomy, we have selected core tasks for speech, music, and audio. These core tasks are reduced subsets of essential tasks that have been widely used or studied within the research community. Specifically, we include tasks from three popular benchmarks: SUPERB (speech), MARBLE (music), and HEAR (audio). The three benchmarks were originally designed for evaluating encoders, not for an instruction-following universal model, so some modifications are required. We manually crafted instructions for each task, reduced the data size, and reformulated them into the Dynamic-SUPERB format. Since the core task set is much smaller than the full benchmark, researchers can more efficiently evaluate their models across a reasonable range of domains.

### v2

@@ Abstract
Multimodal foundation models, such as Gemini and ChatGPT, have revolutionized human-machine interactions by seamlessly integrating various forms of data. Developing a universal spoken language model that comprehends a wide range of natural language instructions is critical for bridging communication gaps and facilitating more intuitive interactions. However, the absence of a comprehensive evaluation benchmark poses a significant challenge. We present Dynamic-SUPERB Phase-2, an open and evolving benchmark for the comprehensive evaluation of instruction-based universal speech models. Building upon the first generation, this second version incorporates 125 new tasks contributed collaboratively by the global research community, expanding the benchmark to a total of 180 tasks, making it the largest benchmark for speech and audio evaluation. While the first generation of Dynamic-SUPERB was limited to classification tasks, Dynamic-SUPERB Phase-2 broadens its evaluation capabilities by introducing a wide array of novel and diverse tasks, including regression and sequence generation, across speech, music, and environmental audio. Evaluation results show that no model performed well universally. SALMONN-13B excelled in English ASR and Qwen2-Audio-7B-Instruct showed high accuracy in emotion recognition, but current models still require further innovations to handle a broader range of tasks. We open-source all task data and the evaluation pipeline at https://github.com/dynamic-superb/dynamic-superb.
@@ 1 Introduction
Despite significant research in universal speech models, evaluating them effectively and comprehensively remains a major challenge. In NLP, benchmarks for text LLMs include a large number of tasks. CrossFit (Ye et al., 2021) includes 160 tasks, BIG-bench (Srivastava et al., 2022) comprises 204 tasks, and Natural-Instructions (Mishra et al., 2022; Wang et al., 2022) provides 1,616 tasks. For universal speech models, several benchmarks have been proposed but with a fixed and limited set of tasks (typically around a dozen) to evaluate specific capabilities of models (Yang et al., 2021; Dunbar et al., 2022; Maimon et al., 2024). This is insufficient for evaluating a universal model, as we aim to investigate whether models can handle a broader range of tasks beyond fixed benchmarks. There is a strong demand for a benchmark that can evaluate these models across various aspects with a wide range of speech tasks, which led to the creation of Dynamic-SUPERB (Huang et al., 2024a).
@@
Dynamic-SUPERB is the first benchmark for evaluating universal instruction-based speech and audio models. As models advance, we dynamically expand the benchmark by adding new tasks to provide better guidance for researchers in model development. We have designed a well-structured pipeline that harnesses the collective efforts of the community to gather diverse challenging, novel, and creative tasks. The first generation of Dynamic-SUPERB consists of 55 tasks, covering various aspects of speech (e.g., semantics, speakers, etc.), making it the speech benchmark with the most tasks. However, this is not sufficient to pave the way for developing universal speech models, especially given the complexity and richness of the information conveyed by spoken language.
@@
This paper presents the Dynamic-SUPERB Phase-2. We expanded it to 180 tasks, more than double the size of its first iteration, with contributions from the global research community. They cover a broad range of types, and some tasks present novel challenges that have not been explored in any previous research. Besides, previous speech research has typically treated music and environmental audio as background noise to be ignored. However, these sounds contain rich information and share overlapping elements that complement each other. Thus, in Dynamic-SUPERB, we also introduced preliminary music and audio tasks from established music and audio benchmarks (Yuan et al., 2023; Turian et al., 2022). We define the core tasks by reformulating the SUPERB (Yang et al., 2021) (speech), MARBLE (Yuan et al., 2023) (music), and HEAR (Turian et al., 2022) (audio) benchmarks for quick-round research experiments. One challenge of evaluating with such a large-scale benchmark is deriving concrete and useful insights from hundreds of evaluation results. We provide a taxonomy for every task in Dynamic-SUPERB, where tasks are clustered by the specific model capabilities they probe. Researchers can follow this taxonomy to develop or reinforce specific capabilities of the models they build. To our knowledge, Dynamic-SUPERB is the largest benchmark and the first to provide such a detailed task taxonomy in speech processing.
@@
We conducted a comprehensive evaluation of several models using Dynamic-SUPERB Phase-2. To handle the diverse output formats of these models, we propose an automated LLM-based pipeline for general evaluation across tasks. The evaluation results show that these models perform well only on a limited range of tasks. For example, SALMONN-13B performed well on English ASR, while Qwen2-Audio-7B-Instruct achieved high accuracy in emotion recognition. However, they do not generalize well to a much broader range of tasks in speech recognition and paralinguistics. Notably, training on diverse data, even with significant differences in signal-level characteristics, can enhance performance across domains. We observed that spoken language models outperformed music models in certain music tasks. This highlights the potential for developing unified models for speech, music, and general audio. We have open-sourced all materials to support reproducibility and further research. We invite researchers to help expand Dynamic-SUPERB, making it more diverse and comprehensive to advance universal spoken language models.
@@ 2.1 Instruction-following universal speech models
LLMs have shown strong natural language processing abilities and are used in speech, audio, and music applications. Recent frameworks integrate a pre-trained speech encoder with an LLM through fine-tuning techniques such as LoRA (Hu et al., 2021). LTU-AS (Gong et al., 2023) combines Whisper (Radford et al., 2023) with LLaMA (Touvron et al., 2023a), and it is fine-tuned on open-ended speech and audio question-answering. SALMONN adopts a window-level Q-former (Li et al., 2023a) to generate soft prompts fusing representations from the speech and audio encoders. Qwen-Audio (Chu et al., 2023) introduces task-specific tags into Qwen to encourage knowledge sharing and prevent interference across diverse tasks, and it supports multi-turn dialogues for both audio and text inputs. WavLLM (Hu et al., 2024) uses curriculum learning to prevent overfitting on specific tasks while maintaining the LLM’s original capabilities. All these models accept speech, audio, and text as input but output only text, focusing on understanding rather than generation in speech and audio. Several attempts have also used prompts for speech, music, and audio generation (Guo et al., 2023; Agostinelli et al., 2023; Liu et al., 2023a). However, to our knowledge, there is no universal model capable of handling both generation and understanding tasks. Moreover, these models have not been comprehensively evaluated on a benchmark, which hinders fair comparisons between them.
@@ 2.2 Evaluation benchmarks
Several benchmarks have been developed to evaluate speech models. SUPERB (Yang et al., 2021) is the most widely used benchmark for assessing the performance of speech foundation models across various tasks that cover different aspects of speech. On the other hand, SLUE (Shon et al., 2022; 2023) focuses more specifically on spoken language understanding. Yet, they are primarily limited to English. LeBenchmark (Evain et al., 2021; Parcollet et al., 2023) evaluates speech foundation models specifically for French, while IndicSUPERB (Javed et al., 2023) is dedicated to Indian languages. ML-SUPERB (Shi et al., 2023; 2024b) offers ASR and language identification tasks in 100+ languages. Aside from speech, MARBLE (Yuan et al., 2023) and HEAR (Turian et al., 2022) offer platforms for evaluating various music and a wide range of audio tasks. Using these benchmarks, researchers build a specialized model for each task. Thus, the number of tasks is limited due to the growing costs associated with adding more tasks. Conversely, universal models are expected to do various tasks without fine-tuning for each one, allowing users to engage in far more diverse applications and enabling benchmark developers to include more tasks for comprehensive evaluation. In NLP and CV, benchmarks have been developed to evaluate models across a much broader range of tasks (Bitton et al., 2023; Srivastava et al., 2022; Mishra et al., 2022). However, in speech, we lack a benchmark of comparable scale. Dynamic-SUPERB (Huang et al., 2024a) is the first benchmark for instruction-based universal speech models. While its first version includes far more tasks than all other speech benchmarks, they are all classification tasks. AIR-bench (Yang et al., 2024b), on the other hand, includes tasks from speech, music, and audio, and expands beyond classification, although the number of tasks available for evaluating universal models remains limited (19 tasks in foundation track). Thus, there remains a critical need for a comprehensive benchmark that evaluates universal models across a broader range of tasks to fully assess their capabilities. Consequently, we developed Dynamic-SUPERB Phase-2, substantially upgrading the first generation with a detailed taxonomy and establishing community contribution protocols to facilitate continual task integration.
@@ 3.2 Task formulation
As Figure 1c depicts, in Dynamic-SUPERB Phase-2, each task is structured to include: (1) Text instruction: A natural language instruction that guides the model on the task to perform. Each task has multiple different instructions to evaluate the model’s ability to understand instructions. (2) Audio component: At least one audio element, which can be in the input, the output, or both. (3) (Optional) Text component: Text elements (other than instruction) that may serve as inputs or outputs. The number of audio elements in the inputs or outputs may vary depending on the task. For example, in speaker verification, two utterances are used to determine whether they were produced by the same speaker (green blocks in Figure 1c). Besides, text inputs are not always required; for instance, ASR does not involve any text inputs. We use text format for instructions instead of spoken format. Spoken instructions involve varying content, speaker characteristics, and prosody, making them more complex. Text instructions act as an intermediary, bridging text and spoken universal models. These designs ensure consistency and simplify the model’s understanding and processing of diverse tasks while maintaining the benchmark’s extendibility with minimal constraints.
@@ 3.3 Call for tasks
Dynamic-SUPERB fosters community collaboration, encouraging the addition of innovative tasks to remain relevant in this rapidly changing field. In Phase-2, we initiated a call for tasks in March 2024 to invite contributions from the research community. We established an organized and transparent submission process on our GitHub portal, where we organize members who serve as editors to guide contributors. Contributors propose tasks by providing relevant information on GitHub. Each proposal is assigned to an editor who checks for major issues and prevents duplicated efforts. Then, task proposers upload their data to our Huggingface space in specified formats, using only datasets with proper licenses. They complete submissions by opening a pull request on GitHub with the necessary files. Afterward, editors review submissions on a rolling basis, offering suggestions for refinement rather than immediate acceptance or rejection. After iterative improvements, accepted tasks are merged into the repository. Between March and July 2024, we received more than 140 task proposals and accepted over 120 new tasks. Tasks still under review are not included here but will be featured in the next phase. For details, please refer to Appendix F.
@@ 3.4 Task taxonomy
Figures 2 shows the high-level task taxonomy in Dynamic-SUPERB Phase-2. Due to the space limitation, we only show some representative tasks. Each leaf node includes at least one task and may be further categorized into more fine-grained sub-domains, which are not shown here due to space constraints. For example, within ‘Speaker & Language/Speaker’, we have two categories: ‘Speaker Characteristics’ and ‘Speaker Identification’, each containing several tasks. Please refer to Appendix B for the complete task list. We first categorize tasks into two primary fields: (I) speech and (II) music & audio, which are generally distinguished by the source of the sound. For instance, speech is produced by human vocal cords, music is created using instruments, and audio includes sounds generated by other creatures, materials, or natural phenomena. We then split each field into several domains based on the key attributes and challenges that the tasks within them present.
@@ 3.4.1 Speech
As Figure 2(a) shows, there are 8 domains within the speech field. Speech Recognition focuses on converting spoken language into text. This includes tackling various challenges like multilingual and code-switching ASR, as well as spontaneous ASR which is very different from audiobook-style ones. It also contains specialized tasks such as command recognition and keyword spotting. Speaker and Language addresses the analysis of speaker characteristics and languages, covering tasks such as speaker verification, diarization, and language identification. Spoken Language Understanding deals with understanding and analysis of the content and semantics of spoken language. It covers tasks like sentiment analysis and speech translation. Phonetics, Phonology, and Prosody focuses on the sound structure of speech, including phoneme recognition, pronunciation evaluation, and prosodic features like stress and accent classification. Paralinguistics explores non-verbal aspects of speech, such as emotion recognition and vocal event detection, which can capture nuances like screaming or coughing. Speech Enhancement aims to improve speech quality by detecting and mitigating noise, reverberation, and other degradations, but currently, we only have understanding tasks. Speech, Voice, and Hearing Disorders is dedicated to identifying and classifying disorders such as stuttering and so on. Safety and Security focuses on detecting synthetic or manipulated speech, addressing tasks like spoof detection and recognizing deepfake voices.
@@ 3.4.3 Core tasks
While our task taxonomy provides a comprehensive, hierarchical, and systematic approach for benchmark users to easily get started with Dynamic-SUPERB Phase-2, evaluating all tasks with limited resources remains a challenge. To address this, alongside the task taxonomy, we have selected core tasks for speech, music, and audio. These core tasks are reduced subsets of essential tasks that have been widely used or studied within the research community. We include tasks from three popular benchmarks: SUPERB (speech), MARBLE (music), and HEAR (audio). The three benchmarks were originally designed for evaluating encoders, not for an instruction-following universal model, so modifications are required. We crafted instructions for each task, reduced the data size, and reformulated them into the Dynamic-SUPERB format. Besides, we replaced the datasets in some tasks to address licensing issues. Since the core task set is much smaller than the full benchmark, researchers can more efficiently evaluate their models across a reasonable range of domains.